# Applicability criteria of proper charge neutrality and special relativistic MHD models extended by two-fluid effects

### v1

@@ Abstract
The applicability of relativistic magnetohydrodynamics (RMHD) and its generalization to two-fluid (including the Hall and inertial effects) is systematically investigated by using the method of dominant balance in the two-fluid equations. Although proper charge neutrality or quasi-neutrality is the key assumption for all MHD models, this condition is difficult to be met when both relativistic and inertial effects are taken into account. The range of application for each MHD model is illustrated in the space of dimensionless scale parameters. Moreover, the number of field variables of relativistic Hall MHD (RHMHD) is shown to be greater than that of RMHD and Hall MHD. Nevertheless, the RHMHD equations may be solved at a lower computational cost than RMHD, since root-finding algorithm, which is the most time-consuming part of the RMHD code, is no longer required to compute the primitive variables.
@@ I Introduction
Magnetized plasmas subject to relativistic effects are common in various high-energy celestial bodies, such as pulsar, black hole magnetosphere, corona of accretion disk, jet from active galactic nucleus and gamma-ray burst. Relativistic magnetohydrodynamics (RMHD) has been used in theoretical and numerical studies as a model to analyze the macroscopic motion of relativistic magnetized plasmas. Since RMHD ignores the microscopic scales of such as inertial length and gyro-radius, it is well known that the magnetic field is frozen in the plasma motion in the collisionless limit. Therefore, RMHD is inappropriate for dealing with magnetic reconnection Uzdensky (2011); Hoshino and Lyubarsky (2012) at least in the microscopic region where magnetic field lines reconnect. In many cases, magnetic reconnection is a key process in which magnetic energy is efficiently converted into kinetic and thermal energy. In addition, RMHD becomes invalid in the limit of low plasma density or weak magnetic field. The Vlasov-Maxwell equations, on the other hand, are based on first principles and have been solved by Particle-In-Cell (PIC) simulations in recent years. However, this direct approach is the most computationally expensive and these kinetic models are difficult to solve analytically. Thus, there is a demand for intermediate models which bridge the gap between RMHD and kinetic ones. In this study, we focus on extended RMHD models that include the two-fluid effects, which are expected to be more widely applicable than RMHD while maintaining a moderate computational cost.
@@
In non-relativistic electron-ion plasmas, a model Loeb (1956); Lüst (1959) including the two-fluid effects (i.e., the Hall effect and the electron inertia effect) is called extended MHD (XMHD) in the recent literature Kimura and Morrison (2014). The XMHD equations are derived from the two-fluid equations by imposing the quasi-neutrality (QN) condition, which approximately eliminates microscopic motions such as plasma oscillation and cyclotron oscillation. XMHD is also shown to have a Hamiltonian structure which conserves canonical vorticities (instead of magnetic flux) Keramidas Charidakos et al. (2014); Abdelhamid, Kawazura, and Yoshida (2015); Lingam, Miloshevich, and Morrison (2016); Hirota (2021). Due to the electron-inertia effect, magnetic reconnection can occur even in the collisionless limit Dungey (1953); Speiser (1970); Ottaviani and Porcelli (1993). Furthermore, the Hall effect is well-known for significantly enhancing the reconnection speed, according to the Global Environment Modeling (GEM) Reconnection Challenge Birn et al. (2001). Since the electron-inertia effect manifests itself on an even smaller scale than the Hall effect, Hall MHD Lighthill (1960) is often used as well, neglecting only the electron-inertia effect. In the case of electron-positron plasmas, the Hall effect vanishes and only the inertial effect remains, so this model is called inertial MHD (IMHD) Kimura and Morrison (2014).
@@
It is natural to assume that there are also some MHD models that include the relativistic effects alongside the two-fluid effects. Such an extension of RMHD in electron-positron plasma was explored early on the literature Lichnerowicz (1967); Anile (2005). Additionally, the extension of generalized Ohm’s law was attempted and applied to pulsar magnetosphereArdavan (1976) [^1] Generalized relativistic Ohm’s law is also proposed by Pegoraro (2015) in a different way.. Koide Koide (2009a) derived a generalized RMHD model from the relativistic two-fluid equations by imposing the proper charge neutrality (PCN) condition, which will be referred to as relativistic extended MHD (RXMHD) in this paper. A variational principle of RXMHD was later proposed by Kawazura et al. Kawazura, Miloshevich, and Morrison (2017). The general relativistic version of RXMHD was presented by KoideKoide (2009b) and by Comisso and Asenjo Comisso and Asenjo (2018) using a covariant form. RXMHD was applied to relativistic collisionless magnetic reconnection Comisso and Asenjo (2014). Relativistic Hall MHD (RHMHD) is similarly obtained by neglecting electron-inertia, and its properties have been studied by Kawazura Kawazura (2017, 2023). However, for the QN condition to hold in non-relativistic MHD, the flow velocity must be sufficiently slower than the speed of light. Moreover, as will be clarified in this paper, the PCN condition in RMHD actually holds in the limit of neglecting the two-fluid effects. Therefore, hybrid models which include both the relativistic and two-fluid effects may violate the charge neutrality condition, requiring careful consideration of the applicability of RXMHD (and RHMHD). In fact, all models bearing the name "MHD" assume either QN or PCN a priori. Once these neutrality conditions (i.e., single-fluid approximation) fail, we should solve the two-fluid equations or kinetic models directly.
@@ II.1 Two-fluid equations
In this study, we start with the special-relativistic fluid equations for both positively and negatively charged gases, where dissipation due to collision is neglected for simplicity. The equations of motion, the continuity equations and Maxwell’s equations are written as
@@ II.2 Transformation into MHD variables
Let us rewrite Eq. (3) and Eq. (4) in terms of MHD variables without any approximation. It is insightful to introduce four-dimensional center-of-mass flux (divided by $c$) by
@@
where $m_{\pm}$ is the mass of particle, and four-dimensional current (with the same dimension as $Q^{\mu}$) by
@@ III.1 Normalization
Because this balance is merely a relation among scale parameters, it should be actually interpreted as $L_{\star}\sim v_{\star}T_{\star}$ or $O(L_{\star})=O(v_{\star}T_{\star})$. But, the equality “$=$” will be used in this paper to reduce the number of the scale parameters by imposing this balance.
Next, consider the Poisson equation (11) which is normalized to
@@
We assume that there is no externally-applied electrostatic potential (e.g., $\phi\rightarrow 0$ at infinity). Then, $\hat{\phi}$ is generated only by the charge density of plasma itself via this equation, and it is natural to assume the balance between the left and right hand sides,
@@
Since the two balances (29) and (31) are assumed among the eight representative scales, let us define 5 dimensionless parameters for later use as follows
@@ III.2 Imposition of MHD balance
Since we are considering magnetized plasmas, this frequency is supposed to be much faster than the time scale of flow dynamics,
$\frac{1/\omega_{c\star}}{T_{\star}}=\frac{\beta_{\star}\epsilon}{\sqrt{\sigma}}\ll 1.$ (54)
The balance (52) indicates that $\epsilon_{m}$ is indeed small and of the same order as $\omega_{c\star}^{-1}/T_{\star}$. It is well-known that the two-fluid equations generally encompass ion’s and electron’s cyclotron motions. By taking the limit of $\epsilon_{m}\rightarrow 0$ while keeping the balance (52), we can eliminate these fast motions from the flow dynamics. We also remark that the $\bm{v}\times\bm{B}$ term (which is $\hat{Q}^{j}\hat{F}^{i}{}_{j}$) in Ohm’s law (III.1) becomes of order $1$ due to (52).
@@
using the balance (52). In the limit of $\epsilon_{m}\rightarrow 0$ or $\sigma\rightarrow 0$, only the electrostatic force term (^$\hat{\partial}_{i}\hat{\phi}$) gets too large $1/(\sigma\epsilon_{m})\rightarrow\infty$ to balance with other terms. This implies that the existence of very fast plasma oscillation breaks down the assumed balance totally. To maintain the balance consistently, the charge separation $\alpha$ must be small enough that all terms in (55) and (56) are equal or less than the order $1$, which requires $\alpha\leq\sqrt{\sigma}\epsilon_{m}$, $\alpha\leq\sigma\epsilon_{m}$ and $\alpha\leq\epsilon_{m}$. To consider the most general situation satisfying all of them, we assume
@@
Since $\epsilon_{\sigma}\leq\sqrt{\sigma}$ holds mathematically (see Fig. 1), the balance (57) leads to
@@ IV.1 Vacuum limit $\beta_{\star}^{2}/\sigma\rightarrow 0$
The limit $\beta_{\star}^{2}/\sigma\rightarrow 0$ corresponds to vacuum state since the right hand side of the Ampere-Maxwell equation (68) vanishes. This limit may be uninteresting because the plasma current is too small to disturb the vacuum magnetic field. For example, the limit of large magnetization parameter $\sigma\rightarrow\infty$ inevitably results in this vacuum state (due to $\beta_{\star}\leq 1$). According to the dispersion relation of RMHD, the relativistic Alfvén velocity $v_{A\star}$ is well-known as
@@ V Range of application
$\epsilon_{m}=\frac{\beta_{\star}}{\sqrt{\sigma}}\epsilon=\frac{v_{\star}}{L_{\star}\omega_{c\star}}\quad\mbox{with}\quad\frac{\beta_{\star}}{\sqrt{\sigma}}\leq 1..$ (85)
@@ VI Remarks on RHMHD
$-\frac{1}{\sigma}(\epsilon_{\sigma}\tilde{n}\bm{E}+\bm{J}\times\bm{B}).$ (100)
This evolution equation for $\bm{E}\times\bm{B}$ can be solved instead of (95) because the electric field has only the perpendicular component and can be reproduced by $\bm{E}=\bm{B}\times(\bm{E}\times\bm{B})/|\bm{B}|^{2}$. Moreover, we can easily eliminate $\epsilon_{\sigma}\tilde{n}\bm{E}+\bm{J}\times\bm{B}$ in (94) and (100) by using Ohm’s law (97). Then, the right hand sides no longer include $\bm{J}$. Therefore, the RHMHD equations are regarded as a dynamical system of 10 fields $(n,\bm{M}_{\rm P},\bm{M}_{\rm EM},\bm{B})$, where $\bm{M}_{\rm P}=n\gamma\bm{v}$ and $\bm{M}_{\rm EM}=\bm{E}\times\bm{B}$, satisfying two constraints $\bm{M}_{\rm EM}\cdot\bm{B}=0$ and $\nabla\cdot\bm{B}=0$. The right hand sides of (93), (94), (96) and (100) are explicitly written in terms of $(n,\bm{M}_{\rm P},\bm{M}_{\rm EM},\bm{B})$, using
@@
In this way, the RHMHD equations are numerically solvable.
@@
where $\bm{M}_{\rm tot}:=n\gamma\bm{v}+\sigma\bm{E}\times\bm{B}$ is the total momentum of plasma and electromagnetic field. According to Ohm’s law, $\bm{E}$ is always replaced by $-\bm{v}\times\bm{B}$. Thus, the RMHD equations are a dynamical system of 7 fields $(n,\bm{M}_{\rm tot},\bm{B})$ with a constraint $\nabla\cdot\bm{B}=0$. However, to calculate the right hand sides of (93), (96) and (103), we need to write $\bm{v}$ in terms of $(n,\bm{M}_{\rm tot},\bm{B})$. It is well known that this is not analytically feasible and requires the use of a root-finding algorithm (such as the Newton-Raphson method). For hot plasma, the evolution equation for the total energy $E_{\rm tot}$ is also solved simultaneously, and the reconstruction of the primitive variables $(n,\bm{v},p,\bm{B})$ from the time-evolving ones $(n,\bm{M}_{\rm tot},E_{\rm tot},\bm{B})$ is one of the most computationally expensive part of RMHD simulation.
In the presence of the Hall effect, we can avoid using root-finding algorithm and the time-marching algorithm becomes straightforward while the number of field variables increases from 7 to 10. The RHMHD equations are possibly solved at a lower cost than the RMHD equations.
@@
The case of $\beta_{\star}^{2}/\sigma\ll 1$ is often uninteresting because it is almost vacuum (i.e., the kinetic energy is much smaller than the energy of externally-applied magnetic field). On the other hand, the inequality $\beta_{\star}^{2}/\sigma\leq 1$ indicates that the maximum velocity scale should be $\beta_{\star}=\sqrt{\epsilon_{\sigma}}$ which is understood as relativistic Alfvén ordering ($v_{\star}\simeq v_{A\star}$). Interesting MHD phenomena are expected in this velocity scale. By focusing on this velocity scale, the number of the nondimensional parameters is further reduced to two $(\epsilon,\sigma)$ which are related to the scales of number density $n_{\star}$, magnetic field $B_{\star}$ and length $L_{\star}$. We have illustrated the applicable ranges of the various MHD models in terms of these scales. For a low density case or in a small scale, it is shown that the charge neutrality condition is violated at an intermediate strength of magnetic field around $\sigma\sim 1$.
We have also summarized the number of field variables for each generalized MHD model. The RHMHD model is shown to be a dynamical system of 10 fields $(n,\bm{v},\bm{B},\bm{E})$ satisfying two constraints ($\nabla\cdot\bm{B}=0$ and $\bm{E}\cdot\bm{B}=0$). This number 10 is different from 7 of the other non-relativistic MHD models and 13 of the original two-fluid model. Moreover, the RHMHD equations describe the time marching of variables $(n,n\gamma\bm{v},\bm{E}\times\bm{B},\bm{B})$ and the primitive variables $(n,\bm{v},\bm{B},\bm{E})$ can be written explicitly by them. Since the RMHD equations requires the root-finding algorithm to calculate $(n,\bm{v},\bm{B})$ from $(n,n\gamma\bm{v}+\sigma\bm{E}\times\bm{B},\bm{B})$, the RHMHD model has an advantage in that the time-marching algorithm is simpler and less expensive than RMHD at the expense of increasing the field variables from 7 to 10. Naturally, RHMHD is a higher fidelity model than RMHD since it includes the Hall effect, which is known to be important for magnetic reconnection process in electron-ion plasma. The application of RHMHD is therefore expected to be beneficial both theoretically and numerically for analysing relativistic plasma phenomena.

### v4

@@ Abstract
The applicability of relativistic magnetohydrodynamics (RMHD) and its generalization to two-fluid models (including the Hall and inertial effects) is systematically investigated by using the method of dominant balance in the two-fluid equations. Although proper charge neutrality or quasi-neutrality is the key assumption for all MHD models, this condition is difficult to be met when both relativistic and inertial effects are taken into account. The range of application for each MHD model is illustrated in the space of dimensionless scale parameters. Moreover, the number of field variables of relativistic Hall MHD (RHMHD) is shown to be greater than that of RMHD and Hall MHD. Nevertheless, the RHMHD equations may be solved at a lower computational cost than RMHD in the limit of cold plasma, since root-finding algorithm, which is the most time-consuming part of the RMHD code, is no longer required to compute the primitive variables.
@@ I Introduction
Magnetized plasmas subject to relativistic effects are common in various high-energy celestial bodies, such as pulsar, black hole magnetosphere, corona of accretion disk, jet from active galactic nucleus and gamma-ray burst. Relativistic magnetohydrodynamics (RMHD) has been used in theoretical and numerical studies as a model to analyze the macroscopic motion of relativistic magnetized plasmas. Since RMHD ignores the microscopic scales of such as inertial length and gyro-radius, it is well known that the magnetic field is frozen in the plasma motion in the zero resistivity limit. Therefore, RMHD is inappropriate for dealing with magnetic reconnection Uzdensky (2011); Hoshino and Lyubarsky (2012) at least in the microscopic region where magnetic field lines reconnect. In many cases, magnetic reconnection is a key process in which magnetic energy is efficiently converted into kinetic and thermal energy. In addition, RMHD becomes invalid in the limit of low plasma density or weak magnetic field. The Vlasov-Maxwell equations, on the other hand, are based on first principles and have been solved by Particle-In-Cell (PIC) simulations in recent years. However, this direct approach is the most computationally expensive and these kinetic models are difficult to solve analytically. Thus, there is a demand for intermediate models which bridge the gap between RMHD and kinetic ones. In this study, we focus on extended RMHD models that include the two-fluid effects, which are expected to be more widely applicable than RMHD while maintaining a moderate computational cost.
@@
In non-relativistic electron-ion plasmas, a model Loeb (1956); Lüst (1959) including the two-fluid effects (i.e., the Hall effect and the electron inertia effect) is called extended MHD (XMHD) in the recent literature Kimura and Morrison (2014). The XMHD equations are derived from the two-fluid equations by imposing the quasi-neutrality (QN) condition, which approximately eliminates microscopic motions such as plasma oscillation and cyclotron oscillation. XMHD is also shown to have a Hamiltonian structure which conserves canonical vorticities (instead of magnetic flux) Keramidas Charidakos et al. (2014); Abdelhamid, Kawazura, and Yoshida (2015); Lingam, Miloshevich, and Morrison (2016); Hirota (2021). Due to the electron-inertia effect, magnetic reconnection can occur even in the zero resistivity limit Dungey (1953); Speiser (1970); Ottaviani and Porcelli (1993).
Furthermore, the Hall effect is well-known for significantly enhancing the reconnection speed, according to the Global Environment Modeling (GEM) Reconnection Challenge Birn et al. (2001). Since the electron-inertia effect manifests itself on an even smaller scale than the Hall effect, Hall MHD Lighthill (1960) is often used as well, neglecting only the electron-inertia effect. In the case of electron-positron plasmas, the Hall effect vanishes and only the inertial effect remains, so this model is called inertial MHD (IMHD) Kimura and Morrison (2014).
@@
It is natural to assume that there are also some MHD models that include the relativistic effects alongside the two-fluid effects. Such an extension of RMHD in electron-positron plasma was explored early on the literature Lichnerowicz (1967); Anile (2005). Additionally, the extension of generalized Ohm’s law was attempted and applied to pulsar magnetosphereArdavan (1976) [^1] Generalized relativistic Ohm’s law is also proposed by Pegoraro (2015) in a different way.. Koide Koide (2009a) derived a generalized RMHD model from the relativistic two-fluid equations by imposing the proper charge neutrality (PCN) condition, which will be referred to as relativistic extended MHD (RXMHD) in this paper. A variational principle of RXMHD was later proposed by Kawazura et al. Kawazura, Miloshevich, and Morrison (2017). The general relativistic version of RXMHD was presented by KoideKoide (2009b) and by Comisso and Asenjo Comisso and Asenjo (2018) using a covariant form. RXMHD was applied to relativistic collisionless magnetic reconnection Comisso and Asenjo (2014). Relativistic Hall MHD (RHMHD) is similarly obtained by neglecting electron-inertia, and its properties have been studied by Kawazura Kawazura (2017, 2023). However, for the QN condition to hold in non-relativistic MHD, the flow velocity must be sufficiently slower than the speed of light. Moreover, as will be clarified in this paper, the PCN condition in RMHD actually holds in the limit of neglecting the two-fluid effects. Therefore, hybrid models which include both the relativistic and two-fluid effects may violate the charge neutrality condition, requiring careful consideration of the applicability of RXMHD (and RHMHD). In fact, all models bearing the name "MHD" assume either QN or PCN a priori. Once these neutrality conditions (i.e., single-fluid approximation) fail, we should solve the two-fluid equations or kinetic models directly. For example, the two-fluid model for relativistic electron-positron plasmas  Barkov et al. has been solved numerically and applied to the simulations of magnetic reconnection Zenitani, Hesse, and Klimas .
@@ II.1 Two-fluid equations
In this study, we start with the special-relativistic fluid equations for both positively and negatively charged gases, where dissipation due to collision is neglected for simplicity. Namely, we ignore resistivity and viscosity. The equations of motion, the continuity equations and Maxwell’s equations are written as
@@ II.2 Transformation into MHD variables
Let us rewrite Eq. (3) and Eq. (4) in terms of MHD variables without any approximation. For this purpose, we define the mass-weighted average of number density $n$, the center-of-mass velocity $v^{i}$, the number density difference $\tilde{n}$, and the velocity $u^{i}$ for the electric current as follows.
$n$ ::=:= $\frac{m_{+}n_{+}+m_{-}n_{-}}{m_{+}+m_{-}},\hskip 10.0ptv^{i}:=\frac{m_{+}n_{+}v^{i}_{+}+m_{-}n_{-}v^{i}_{-}}{(m_{+}+m_{-})n},$
$\tilde{n}$ ::=:= $n_{+}-n_{-},\hskip 10.0ptu^{i}:=\frac{n_{+}v^{i}_{+}-n_{-}v^{i}_{-}}{n},$ (13)
where $m_{\pm}$ is the mass of particle. Moreover, these variables are associated with four-dimensional center-of-mass flux (divided by $c$),
@@
and four-dimensional current (with the same dimension as $Q^{\mu}$),
@@ III.1 Normalization
Since this balance is merely a relation among scale parameters, it should be actually interpreted as $L_{\star}\sim v_{\star}T_{\star}$ or $O(L_{\star})=O(v_{\star}T_{\star})$. However, in this paper, the equality “$=$” will be used to reduce the number of the scale parameters by imposing this balance.
Next, we consider the Poisson equation (11) which is normalized to
@@
and assume that there is no externally-applied electrostatic potential (e.g., $\phi\rightarrow 0$ at infinity). Then, $\hat{\phi}$ is generated only by the charge density of plasma itself via this equation, and it is natural to assume the balance between the left and right hand sides,
@@
Since the two balances 1 and 2 are assumed among the eight representative scales, let us define 5 dimensionless parameters for later use as follows
@@ III.2 Imposition of MHD balance
Then, the balance 5 indicates that
$\epsilon_{m}=\frac{1/\omega_{c\star}}{T_{\star}}$ (55)
is small if this frequency is faster than the time scale $T_{\star}$ of flow dynamics. It is well known that the two-fluid equations generally encompass ion’s and electron’s cyclotron motions. By taking the limit of $\epsilon_{m}\rightarrow 0$ while keeping the balance 5, we can eliminate these fast motions from the flow dynamics. We also remark that the $\bm{v}\times\bm{B}$ term (which is $\hat{Q}^{j}\hat{F}^{i}{}_{j}$) in Ohm’s law (III.1) becomes of order $1$ due to (53).
Although the balance 5 is assumed in this work, we do not claim here that Hall or extended MHD must always satisfy $\epsilon_{m}=\omega_{c\star}^{-1}/T_{\star}\ll 1$. In electron-ion plasmas, $\omega_{c\star}$ is approximately the ion cyclotron frequency $\omega_{ci}$, and the frequency of the whistler wave is known to be higher than $\omega_{ci}$, i.e. $\epsilon_{m}>1$. As will be shown later in Sec. IV and V, the electron inertia effect ($\epsilon_{I}^{2}$) can be neglected due to the smallness of the mass ratio $m_{-}/m_{+}$. If we take the limit of $m_{-}/m_{+}\rightarrow 0$, the Hall and extended MHD would be applicable even if $\epsilon_{m}>1$.
@@
using the balance 5. In the limit of $\epsilon_{m}\rightarrow 0$ or $\sigma\rightarrow 0$, only the electrostatic force term (^$\hat{\partial}_{i}\hat{\phi}$) gets too large $1/(\sigma\epsilon_{m})\rightarrow\infty$ to balance with other terms. This implies that the existence of very fast plasma oscillation breaks down the assumed balance totally. To maintain the balance consistently, the charge separation $\alpha$ must be small enough that all terms in (56) and (57) are equal or less than the order $1$, which requires $\alpha\leq\sqrt{\sigma}\epsilon_{m}$, $\alpha\leq\sigma\epsilon_{m}$ and $\alpha\leq\epsilon_{m}$. To consider the most general situation satisfying all of them, we assume
@@
Since $\epsilon_{\sigma}\leq\sqrt{\sigma}$ holds mathematically (see Fig. 1), the balance 6 leads to
@@ IV.1 Vacuum limit $\beta_{\star}^{2}/\sigma\rightarrow 0$
The limit $\beta_{\star}^{2}/\sigma\rightarrow 0$ corresponds to vacuum state since the right hand side of the Ampere-Maxwell equation (69) vanishes. This limit may be uninteresting because the plasma current is too small to disturb the vacuum magnetic field. For example, the limit of large magnetization parameter $\sigma\rightarrow\infty$ inevitably results in this vacuum state (due to $\beta_{\star}\leq 1$)
As another high magnetization limit $\sigma\rightarrow\infty$, the force-free approximation is widely used Komissarov . When the current is divided into parallel ($J_{\parallel}$) and perpendicular ($\bm{J}_{\perp}$) components to the magnetic field, $J_{\parallel}$ remains finite while $\bm{J}_{\perp}=0$ (i.e., $\bm{J}\times\bm{B}=0$) in the force-free state. In this study, we set $n_{\star}u_{\star}$ as the single representative scale of the current, and no distinction is made between the parallel and perpendicular components. To obtain the force-free state, only the parallel component should diverge ($u_{\parallel\star}\rightarrow\infty$) to keep the parallel current finite in the vacuum limit $n_{\star}\rightarrow 0$. Although the force-free model will be reproduced later in Sec. VI, the existence of the "extended" force-free model is non-trivial and will be the subject of future work.
According to the dispersion relation of RMHD, the relativistic Alfvén velocity $v_{A\star}$ is well-known as
@@ IV.3 Non-relativistic limit $\sigma\rightarrow 0$
The equation (80) is the generalized Ohm’s law. Recall that we neglected resistivity and viscosity at the beginning, assuming that they are not dominant. The inclusion of these terms with large Lundquist and Reynolds numbers does not affect the balances we have discussed so far.
@@ V Range of application
$\epsilon_{m}=\frac{\beta_{\star}}{\sqrt{\sigma}}\epsilon=\frac{v_{\star}}{L_{\star}\omega_{c\star}}\quad\mbox{with}\quad\frac{\beta_{\star}}{\sqrt{\sigma}}\leq 1.$ (86)
@@ VI Remarks on RHMHD
$-\frac{1}{\sigma}(\epsilon_{\sigma}\tilde{n}\bm{E}+\bm{J}\times\bm{B}),$ (101)
$\partial_{t}(\bm{E}\cdot\bm{B})=$ $\frac{1}{\beta_{\star}^{2}}\bm{B}\cdot(\nabla\times\bm{B})-\bm{E}\cdot(\nabla\times\bm{E})-\frac{\bm{J}\cdot\bm{B}}{\sigma}.$ (102)
To preserve the constraint $\bm{E}\cdot\bm{B}=0$, the current parallel to the magnetic field ($\bm{J}\cdot\bm{B}$) is determined such that the right hand side of (102) becomes always zero.
$\frac{\bm{J}\cdot\bm{B}}{\sigma}=\frac{1}{\beta_{\star}^{2}}\bm{B}\cdot(\nabla\times\bm{B})-\bm{E}\cdot(\nabla\times\bm{E}).$ (103)
On the other hand, the evolution equation (101) for $\bm{E}\times\bm{B}$ can be solved instead of (96) because the electric field has only the perpendicular component and can be reproduced by $\bm{E}=\bm{B}\times(\bm{E}\times\bm{B})/|\bm{B}|^{2}$. Moreover, we can easily eliminate $\epsilon_{\sigma}\tilde{n}\bm{E}+\bm{J}\times\bm{B}$ in (95) and (101) by using Ohm’s law (98). Then, the right hand sides no longer include $\bm{J}$. Therefore, the cold RHMHD equations are regarded as a dynamical system of 10 fields $(n,\bm{M}_{\rm P},\bm{M}_{\rm EM},\bm{B})$, where $\bm{M}_{\rm P}=n\gamma\bm{v}$ and $\bm{M}_{\rm EM}=\bm{E}\times\bm{B}$, satisfying two constraints $\bm{M}_{\rm EM}\cdot\bm{B}=0$ and $\nabla\cdot\bm{B}=0$. The right hand sides of (94), (95), (97) and (101) are explicitly written in terms of $(n,\bm{M}_{\rm P},\bm{M}_{\rm EM},\bm{B})$, using
@@
In this way, the RHMHD equations are numerically solvable in the cold plasma limit.
We also note that the force-free state can be obtained in the high magnetization limit $\sigma\rightarrow\infty$, where only (97) and (101) are solved by neglecting the last term of (101). Although $\bm{J}$ is already eliminated from these equations, it must exist at orders $\bm{J}\times\bm{B}=O(1)$ and $\bm{J}\cdot\bm{B}=O(\sigma)$. Our normalization $\bm{J}=O(1)$ is therefore violated in the parallel component and needs to be modified to treat this model.
@@
where $\bm{M}_{\rm tot}:=n\gamma\bm{v}+\sigma\bm{E}\times\bm{B}$ is the total momentum of plasma and electromagnetic field. According to Ohm’s law, $\bm{E}$ is always replaced by $-\bm{v}\times\bm{B}$. Thus, the RMHD equations are a dynamical system of 7 fields $(n,\bm{M}_{\rm tot},\bm{B})$ with a constraint $\nabla\cdot\bm{B}=0$. However, to calculate the right hand sides of (94), (97) and (106), we need to write $\bm{v}$ in terms of $(n,\bm{M}_{\rm tot},\bm{B})$. It is well known that this is not analytically feasible and requires the use of a root-finding algorithm (such as the Newton-Raphson method), which is one of the most computationally expensive part of RMHD simulation.
For the case of hot plasma, the momentum equation (95) is rewritten as
$\partial_{t}(h\gamma^{2}\bm{v})=-\nabla p-\nabla\cdot(h\gamma^{2}\bm{v}\bm{v})+\epsilon_{\sigma}\tilde{n}\bm{E}+\bm{J}\times\bm{B},$ (107)
including the enthalpy density $h$ and pressure $p$. The evolution equation for the total energy $E_{\rm tot}$ should be solved additionally when the barotropic condition is not satisfied, and the primitive variables $({N},\bm{v},p,\bm{B})$ should be reconstructed from the time-evolving variables. Since $M_{p}=h\gamma^{2}v$ depends on $N=n/\gamma$ even in the barotropic case $h(N)$, the formula (104) no longer determines $\bm{v}$ explicitly. Unfortunately, a root-finding algorithm is again necessary for RHMHD.
Therefore, in the presence of the Hall effect and the cold plasma limit, we can avoid using root-finding algorithm and the time-marching algorithm becomes straightforward while the number of field variables increases from 7 to 10. The cold RHMHD equations are possibly solved at a lower cost than the RMHD equations.
@@
The case of $\beta_{\star}^{2}/\sigma\ll 1$ is often uninteresting because it is almost vacuum (i.e., the kinetic energy is much smaller than the energy of externally-applied magnetic field). In this paper, we set aside the force-free state for simplicity, which allows for only the existence of finite parallel current even in the vacuum limit. On the other hand, the inequality $\beta_{\star}^{2}/\sigma\leq 1$ indicates that the maximum velocity scale should be $\beta_{\star}=\sqrt{\epsilon_{\sigma}}$ which is understood as relativistic Alfvén ordering ($v_{\star}\simeq v_{A\star}$). Interesting MHD phenomena are expected in this velocity scale. By focusing on this velocity scale, the number of the nondimensional parameters is further reduced to two $(\epsilon,\sigma)$ which are related to the scales of number density $n_{\star}$, magnetic field $B_{\star}$ and length $L_{\star}$. We have illustrated the applicable ranges of the various MHD models in terms of these scales. For a low density case or in a small scale, it is shown that the charge neutrality condition is violated at an intermediate strength of magnetic field around $\sigma\sim 1$.
We have also summarized the number of field variables for each generalized MHD model. The RHMHD model is shown to be a dynamical system of 10 fields $(n,\bm{v},\bm{B},\bm{E})$ satisfying two constraints ($\nabla\cdot\bm{B}=0$ and $\bm{E}\cdot\bm{B}=0$). This number 10 is different from 7 of the other non-relativistic MHD models and 13 of the original two-fluid model. Moreover, the cold RHMHD equations describe the time marching of variables $(n,n\gamma\bm{v},\bm{E}\times\bm{B},\bm{B})$ and the primitive variables $(n,\bm{v},\bm{B},\bm{E})$ can be written explicitly by them. Since the cold RMHD equations requires the root-finding algorithm to calculate $(n,\bm{v},\bm{B})$ from $(n,n\gamma\bm{v}+\sigma\bm{E}\times\bm{B},\bm{B})$, the cold RHMHD model has an advantage in that the time-marching algorithm is simpler and less expensive than RMHD at the expense of increasing the field variables from 7 to 10. Naturally, RHMHD is a higher fidelity model than RMHD since it includes the Hall effect, which is known to be important for magnetic reconnection process in electron-ion plasma. The application of RHMHD is therefore expected to be beneficial both theoretically and numerically for analyzing relativistic plasma phenomena.
Acknowledgements.
We thank Y. Kawazura and K. Toma for helpful discussion. This work was supported by JST SPRING, Grant Number JPMJSP2114, a Scholarship of Tohoku University, Division for Interdisciplinary Advanced Research and Education, and IFS Graduate Student Overseas Presentation Award.